OpenStudy (anonymous):

The sum of two numbers is 6. The difference of their squares is 12. What is the positive difference of the two numbers?

OpenStudy (calculusxy):

Let's start by having addends of 2 and 4. We know that 2+ 4=6. When you square 2, you get 4 (2*2= 4) and when you square 4 you get 16 (4*4=16). And so when you subtract 16-4=12.

OpenStudy (mathstudent55):

Let the numbers be x and y. "The sum of two numbers is 6.": x + y = 6 "The difference of their squares is 12.": x^2 - y^2 = 12 Solve the system of equations: x + y = 6 x^2 - y^2 = 12